OpenStudy (anonymous):

Mr. Jones has taken a survey of college students and found that 1 out of 7 students are liberal arts majors. If a college has 11,000 students, what is the best estimate of the number of students who are liberal arts majors

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

1572

OpenStudy (anonymous):

whos owlfrred lol

OpenStudy (anonymous):

1 out of 7 students are liberal arts majors so the ratio is 1/7 if a college has 11000 students -> 1/7 = x/11000 x=11000/7=1571.4 , x approximates 1571 x is the number of students who are liberal arts major

OpenStudy (anonymous):

yeah what metarock said. if you don't round the answer is 1571.4 but i assumed youd round up since you can't have .4 of a person lol..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Okay. Thanks !